OpenStudy (anonymous):

square root x + 4 = -3

OpenStudy (anonymous):

x+4=-3 -4 -4 --------------------- X=-7 \[\sqrt{-7}\]

OpenStudy (anonymous):

I dont think you would really wanna simplify that.

OpenStudy (anonymous):

thank you....so much...

OpenStudy (anonymous):

Np!

OpenStudy (anonymous):

2.64575131 i

OpenStudy (anonymous):

hmm Let me check.

OpenStudy (anonymous):

\[\sqrt{7}\approx2.6458\]

OpenStudy (anonymous):

love you

OpenStudy (anonymous):

:) Ditto!